### v1

# X-GAN: A Generative AI-Powered Unsupervised Model for High-Precision Segmentation of Retinal Main Vessels toward Early Detection of Glaucoma

Abstract
Structural changes in main retinal blood vessels serve as critical biomarkers for the onset and progression of glaucoma. Identifying these vessels is vital for vascular modeling yet highly challenging. This paper proposes X-GAN, a generative AI-powered unsupervised segmentation model designed for extracting main blood vessels from Optical Coherence Tomography Angiography (OCTA) images. The process begins with the Space Colonization Algorithm (SCA) to rapidly generate a skeleton of vessels, featuring their radii. By synergistically integrating generative adversarial networks (GANs) with biostatistical modeling of vessel radii, X-GAN enables a fast reconstruction of both 2D and 3D representations of the vessels. Based on this reconstruction, X-GAN achieves nearly 100% segmentation accuracy without relying on labeled data or high-performance computing resources. Also, to address the Issue, data scarity, we introduce GSS-RetVein, a high-definition mixed 2D and 3D glaucoma retinal dataset. GSS-RetVein provides a rigorous benchmark due to its exceptionally clear capillary structures, introducing controlled noise for testing model robustness. Its 2D images feature sharp capillary boundaries, while its 3D component enhances vascular reconstruction and blood flow prediction, supporting glaucoma progression simulations. Experimental results confirm GSS-RetVein’s superiority in evaluating main vessel segmentation compared to existing datasets. Code and dataset are here: https://github.com/VikiXie/SatMar8.
1 Introduction
Glaucoma is one of the primary causes of irreversible blindness and vision impairment [35, 37, 58]. Early detection is crucial, as the disease often progresses silently, with noticeable symptoms appearing only after significant damage has already occurred [21, 22]. Recent studies have identified changes in choroidal blood vessels, especially microvasculature dropout, as promising new biomarkers for monitoring glaucoma progression [41, 54]. These changes can be assessed using the advanced imaging technology of Optical Coherence Tomography Angiography (OCTA) techniques [30, 23]. However, accurately evaluating choroidal vessel density using OCTA is a complex and nontrivial task.
Existing medical imaging analytics in glaucoma research, such as extraction and segmentation of major retinal blood vessels, predominantly depend on computer vision techniques [29, 27, 31, 28, 23, 37, 58, 57, 38, 36, 35, 19]. In recent years, supervised learning segmentation models (SLSMs) [49, 19, 44, 42, 31, 27, 28, 16, 39, 29, 62, 6, 64, 9] have become particularly prominent in this field. However, SLSMs may not be well-suited for detecting choroidal vessels in OCTA images. Supervised learning approaches rely heavily on labeled data, but manually annotating choroidal blood vessels in OCTA images is extremely challenging with current devices [30, 40]. OCTA images offer high-resolution, depth-resolved visualization of retinal and choroidal microvasculature without the need of contrast dye, making them more effective than traditional fundus imaging for detecting vascular abnormalities [30, 40]. However, segmenting choroidal vessels is difficult due to irregular patterns, intersecting pathways, dense capillary networks [26, 21], and complex capillary structures. Moreover, the presence of major blood vessels, which must be excluded, further complicates the labeling process [21, 22].
To fill the gap, this research introduces a novel unsupervised approach to detecting blood vessels in OCTA images.
Glaucoma researchers have revealed that retinal blood vessels follow biostatistical relationships, meaning that vessel radius variations along branching structures [7, 23]. raising the question of whether segmentation can bypass pixel-based mapping and leverage vessel radius as a key marker. To address this, we propose X-GAN, an unsupervised model that integrates GANs [13] with retinal vessel biostatistics for precise main vessel segmentation from OCTA images, eliminating the need for labeled data or extensive training. To tackle data scarcity, we introduce GSS-RetVein, a benchmark dataset with clear capillary distribution in 2D and 3D, enhancing segmentation, 3D reconstruction, and blood flow prediction.
In summary, the contributions of this paper are three-fold:
  • We introduce X-GAN, a generative AI-powered unsupervised segmentation model tailored for extracting main blood vessels from OCTA images.
  • We propose the GSS-RetVein, a unique and pioneering dataset focused exclusively on retinal vascularity. GSS-RetVein consists of 2D and 3D imaging data, and also preserves intricate capillary details without compromising resolution, ensuring sharp blood vessel boundaries and minimal pixel-induced blurring.
  • Experimental results of X-GAN on GSS-RetVein, OCTA-500 [30] and ROSE [40] for retinal main vessel segmentation have demonstrate that X-GAN outperformed state-of-the-art (SOTA) cureent models, achieving nearly 100% segmentation accuracy. Also, experimental results of different models on GSS-RetVein have shown the superiority of GSS-RetVein.
2 Related Work
2.1 Datasets for Glaucoma
Many datasets of glaucoma focus on computer vision tasks [37, 58, 57, 38, 36, 35, 28, 40, 30, 63, 52, 53]. For inner retinal blood vessels, datasets like LAG [28], OCTA-500 [30] and ROSE [40], provide various-sizeimages (e.g., 3mm, 6mm). However, these images represent only a small portion of the total dataset. Some datasets [63, 52, 53, 37, 58, 57, 38, 36, 35, 28] offer fundus images, and the capillaries are nearly indistinguishable to the naked eye. There is a shortage of high-quality, intuitive data, such as detailed images of inner 3mm×$\times$3mm retinal vessels, and existing datasets often lack sufficient clarity. Consequently, few studies focus on structural changes in major retinal blood vessels in glaucoma.
Figure 1: The architecture of X-GAN. Red indicates that X-GAN does not utilize these things in Segmentor $S$, whereas GAN+SLSMs do.
2.2 Computer Vision for Glaucoma
Even when data is accessible, challenges in labeling and the subjective variability among doctors may introduce biases or hinder optimal model fitting [37, 58, 57, 38, 36, 35, 19]. Generative AI can compensate for data deficiencies and labeling issues, though it often requires human verification [60] and generated data also face challenges in labeling. However, GAN [13, 66, 20, 65, 24, 61] is an unsupervised model which can solve both issues: data scarity and labeling. GAN-based segmentation methods are generally designed in conjunction with a dedicated segmentation network. For the segmentor part $S$, many pre-trained models like CNNs [27, 31, 28], U-Net [49, 64, 19, 44], ViT [10], Mask R-CNN [16] and MedSAM [39] can be used for final segmentation. Most recent studies from CVPR [50, 20, 2, 8], IEEE Transactions on Medical Imaging [23, 61, 11, 24] and other top conferences or journals [59, 3, 65] also follow this way.
2.3 Motivation
For GAN+SLSM [23, 61, 2, 11, 24, 20], or SLSMs [49, 19, 44, 42, 31, 27, 28, 16, 39, 29, 62, 6, 64, 9], pixel-level representations can be subjective, as different doctors may annotate the details of the same object differently [35, 30, 29]. In OCTA imaging, retinal blood vessels exhibit a uniform color change and have a shape closely resembling a path. They can be represented using coordinates along with road width. Space Colonization Algorithm (SCA) [46] just meets this requirement. Rather than generating pixel-wise vessel segmentations, we formulates the vascular network as a structured graph representation, ensuring topological consistency and robust connectivity. The graph consists of the following components, shown in Eq. 1:
$V=\left\{(x_{i},y_{i},z_{i}|r_{i})\right\}_{i=1}^{N}$ (1)
where ($x_{i}$, $y_{i}$, $z_{i}$) are the spatial coordinates of the vessel centerline points, $r_{i}$ represents the local vessel radius and $N$ means the number of nodes. This structured representation is initially generated using the SCA [46, 32, 25], which models vessel growth based on attraction points and bifurcation principles. However, the initial vessel maps exhibit domain discrepancies in contrast and noise characteristics compared to real OCTA images. To mitigate this gap, we employ a tuned GAN based on CycleGAN [66] as a structural refinement module, aligning vessel representations with real OCTA distributions while preserving vascular topology.
We introduce the Depth-First Search (DFS, a nineteenth century algorithm)-based segmentation approach [55], applying DFS directly to structured (coordinates, radius) data before image rendering. Unlike GAN+SLSM [23, 61, 2, 11, 24, 20], which segments generated images mixed with real data, our method intercepts the GAN generation process, ensuring segmentation aligns with biostatistical vessel radii. The DFS threshold, $R_{min}$, is defined as the minimum radius ratio of main vessels, following established biostatistical principles [22, 12, 5, 43, 7].
Also, we introduce GSS-RetVein, a dataset capturing rich details of both main vessels and capillaries. As shown in Fig. 4, its images (2D, Fig. 4(b)) retain clear capillary edges and tips, unlike ROSE [40] (Fig. 4(d)) and OCTA-500 [30] (Fig. 4(c)). The dense capillaries introduce noise, effectively testing model robustness for SLSMs. This detail aids vascular reconstruction, blood flow prediction, and glaucoma progression simulation by enhancing edge and endpoint learning [30, 29, 62]. Inspired by the Harvard glaucoma dataset series [37, 58, 57, 38, 36, 35], we integrate fairness measures to minimize bias and ensure equitable model performance.
Note: given the model’s simplicity (adjusted GAN + DFS) yet crazy performance (nearly 100% segmentation accuracy), we name it ”X” to highlight its exceptional segmentation ability, similar to how Elon Musk rebranded Twitter as X. Thus, we refer to our model as X-GAN.
3 Methodology: X-GAN
The complete framework of X-GAN is illustrated in Fig. 1 and it consists of two parts: Vessel Structure Refinement Module Module and Segmentor.
3.1 Vessel Structure Refinement Module
Given a vessel structure $X_{s}$ via SCA, our generator $G$ learns to transform it into a realistic OCTA-like representation $X_{r}^{\prime}$, expressed as $G(X_{s})$, while $X_{r}^{\prime}$ undergoes style adaptation to align with the contrast and noise characteristics of real OCTA images $X_{r}$. To enhance structural consistency, we modify the original CycleGAN architecture by removing the inverse generator and incorporating a segmentation consistency loss via the segmentor $S$. This modification prevents CycleGAN from altering the vessel topology, ensuring that vessel segmentation remains biologically meaningful, as illustrated in Eq. 2:
$L_{\text{seg}}=\mathbb{E}_{X_{r}}\left[\|S(X_{r}^{\prime})-S(X_{r})\|_{1}\right]$ (2)
where $S(X)$ extracts vessel masks from OCTA images, enforcing structural consistency during transformation. This constraint ensures that CycleGAN adapts contrast while strictly preserving vessel topology. Consequently, our total loss function is formulated as follows:
$L_{\text{GAN}}$ $=\mathbb{E}_{X_{r}}[D(X_{r})]-\mathbb{E}_{X_{s}}[D(G(X_{s}))]$ (3)
$L_{\text{GP}}$ $=\lambda\mathbb{E}_{\hat{X}}\left[\left(\|\nabla_{\hat{X}}D(\hat{X})\|_{2}-1\right)^{2}\right]$ (4)
$L_{\text{total}}$ $=L_{\text{GAN}}+\lambda_{\text{GP}}L_{\text{GP}}+\lambda_{\text{seg}}L_{\text{seg}}$ (5)
where Eq. 3 is the Wasserstein Adversarial Loss, Eq. 4 is the Gradient Penalty for Stability ($\hat{X}$ is the reference vessel representation) and Eq. 5 is the Structural Consistency via Segmentation Loss. Upon training convergence, the generator produces a vessel structure that not only exhibits realistic contrast properties but also preserves anatomical coherence, ensuring fidelity to real OCTA characteristics.
3.2 Segmentor
Rather than utilizing pixel-wise segmentation maps, we directly extract primary vessels from the $(x_{i}^{G},y_{i}^{G},z_{i}^{G}|r_{i}^{G})$ representation using DFS based graph traversal, which effectively isolates large vessels while filtering out capillaries.
First, we construct a graph $G=(V,E)$, where nodes $V$ represent vessel centerline points and edges $E$ connect adjacent vessel points based on local vessel connectivity. A vessel segment $e_{ij}$ between nodes $v_{i}$ and $v_{j}$ (the direction is from $i$ to $j$) is assigned a weight, as shown in Eq. 6:
$w(e_{ij})=r_{i}$ (6)
where larger vessels are prioritized in the traversal process. To extract primary vessels, we apply radius thresholding and DFS traversal. For radius filtering, we define a minimum main vessel radius $R_{min}$ and retain only Eq. 7:
$V_{\text{main}}=\left\{(x_{i},y_{i},z_{i}|r_{i})\mid r_{i}\geq(R_{\min}*r_{max})\right\}$ (7)
For DFS traversal, it consists of three steps: (1) select the optic disc region as the root node; (2) recursively traverse the largest connected component using DFS; (3) stop when no further vessel segments satisfy the radius constraint. When all three steps finished, construct final extracted vessel structure $G_{main}$. This approach effectively retains only the primary vascular structure, removing small capillaries and noise. This method can be summarized as the following Algorithm 1.
Algorithm 1 : Primary Vessel Extraction via DFS
1:Vessel structure $V=\{(x_{i},y_{i},z_{i}|r_{i})\}_{i=1}^{N}$
2:Extracted primary vessel graph $G_{\text{main}}$
3:Step 1: Vessel Graph Construction
4:for all vessel point $(x_{i},y_{i},z_{i}|r_{i})\in V$ do
5:     Identify neighboring vessel points to form edges $E$
6:     Assign edge weights: $w(e_{ij})=r_{i}$
7:end for
8:Step 2: Vessel Filtering
9:Apply radius threshold $R_{\text{min}}$:
10:$V_{\text{main}}=\{(x_{i},y_{i},z_{i}|r_{i})\mid r_{i}\geq(R_{\text{min}}*r_{max})\}$
11:Remove weakly connected components
12:Step 3: DFS for Vessel Traversal
13:Initialize DFS from optic disc region
14:for all vessel point $(x_{i},y_{i},z_{i}|r_{i})\in V_{\text{main}}$ do
15:     Recursively explore vessel network using DFS
16:     Expand along larger vessels
17:end for
18:Step 4: Output Primary Vessel Graph return $G_{\text{main}}$
The key aspect of Algorithm 1 is the threshold $R_{min}$. Instead of the final image, we analyze intermediate outputs $S(X_{r})$, CSV data (Fig. 2(e)), and GAN-derived radius statistics (Fig. 2(d)). For a sample image (Fig. 2(a)), we computed a Mean (85.31) and Standard Deviation (88.67), determining $R_{min}$ as 0.1968 $R_{max}$. Across datasets, values were 0.2011 (GSS-RetVein), 0.2102 (OCTA-500), and 0.1982 (ROSE), standardized to 0.2 for consistency. Ophthalmology studies [22, 12, 5, 43, 7] confirm 18-26% capillary coverage in OCTA images, validating this threshold.
(a) GSS-RetVein
(b) 3D Visualization
(c) Statistical Distribution
(d) Radius Statistics
(e) Vascular Data Format (Part)
Figure 2: 3D visualization of the pixel distribution of one sample, and its Gaussian statistical distribution.
[…]
4 Dataset and Implement
4.1 Dataset
4.1.1 GSS-RetVein
GSS-RetVein is a unique and pioneering dataset focused exclusively on retinal vascularity. The dataset features an exceptionally clear distribution of capillaries. It comprises 550 images, both 2D and 3D, from 250 subjects, ensuring fair representation across gender and race. We collected retinal vascular data for both eyes (OS and OD) of each subject, with a gender split of 48.8% male and 51.2% female. Racial distribution reflects local glaucoma clinic demographics: 80.4% white, 14.6% Black, with Asians and other minority groups comprising the remainder. Additional data includes age (47.1 $\pm$ 24.8 years) and examination dates from 2023 to 2024.
GSS-RetVein’s original images come from the latest Intalight device [^1] https://intalight.com/, offering superior detail compared to OPTOVUE [^2] https://www.visionix.com/ (Fig. 3). As shown in Fig. 3(a) and Fig. 3(c), Intalight images feature clearer vessel boundaries and capillary distributions, enhancing segmentation accuracy. In supervised learning, their minimal pixel variation introduces valuable noise for testing model robustness [62]. For 3D reconstruction and evolution prediction, these high-resolution images aid in vascular modeling and blood flow analysis [29], making Intalight data ideal for both tasks.
(a) Intalight
(b) OPTOVUE
(c) Intalight
(d) OPTOVUE
Figure 3: Comparison of data collected by different devices for the same patient. Among them, Fig. 3(a) and Fig. 3(b) are images of deep angio, and Fig. 3(c) and Fig. 3(d) are images of superficial angio. They all are images of inner retina vessel.
For GSS-RetVein, we eliminate red and green baselines from Intalight device images (Fig. 4(a)) through a multi-step process. First, we apply a dark channel algorithm [15] to remove light artifacts, followed by unsharp masking [51] to enhance high-frequency details. The final images (Fig. 4(b)) undergo manual verification for quality assurance. As shown in Fig. 4, while main vessels remain visible across datasets, magnification reveals that OCTA-500 [30] and ROSE [40] blur capillaries and vessel ends. Even the unprocessed baseline image (Fig. 4(a)) offers superior vessel clarity compared to other datasets (Fig. 4(d), Fig. 4(c)).
(a) Baseline
(b) GSS-RetVein
(c) OCTA-500
(d) ROSE
Figure 4: Comparison of GSS-RetVein images with other datasets.
For the representation of 3D imaging of GSS-RetVein, as shown in Fig. 5, capillaries stand out from the thicker main blood vessels by attaching to them and narrowing at their ends (Fig. 5(c)). The elaborate architecture of blood vessels and the intertwining capillaries contribute to an extremely intricate vascular layout. 3D imaging part of GSS-RetVein is formatted as a JSON file, which includes 3D X, Y, and Z coordinates and the corresponding radius pixel distribution values. For annotation (2D, image format: PNG), our team includes two glaucoma specialists who handle labeling, verification, and manual evaluation, utilizing the Labelme [^3] https://github.com/wkentaro/labelme.
(a) Top View
(b) Front View
(c) Detail
Figure 5: 3D imaging of retinal vessel shown from various angles.
In general, GSS-RetVein is a dataset that combines 2D and 3D images with extremely high image resolution, and is particularly dedicated to retinal vascular segmentation.
Model Dataset
GSS-RetVein OCTA-500 ROSE
6mm 3mm 6mm 3mm 3mm
IoU Dice IoU Dice IoU Dice IoU Dice IoU Dice
U-Net [49] 95.14 97.51 95.23 97.56 96.93 98.44 98.75 99.37 97.94 98.96
U-Net++ [64] 95.68 97.81 96.64 98.29 97.65 98.81 97.96 98.97 97.52 98.75
Attention U-Net [44] 95.78 97.84 96.04 97.98 97.05 98.50 98.61 99.30 97.92 98.95
Mask R-CNN [16] 92.35 96.02 91.35 95.48 92.06 95.87 93.50 96.64 91.23 95.41
YOLOv11-x [18] 95.81 97.86 96.55 98.24 95.93 97.92 97.66 98.82 97.83 98.89
MedSAM [39] 96.38 98.16 96.47 98.20 96.01 97.95 98.94 99.47 97.89 98.93
CauSSL [42] 95.23 97.56 95.47 97.68 96.32 98.04 98.15 99.08 97.80 98.89
UniverSeg [6] 95.30 97.58 95.52 97.70 96.58 98.17 98.42 99.22 97.07 98.01
S2VNet [9] 95.68 97.81 95.36 97.63 96.45 98.11 98.27 99.15 97.92 98.95
Tyche [45] 95.80 97.85 95.19 97.53 96.28 98.02 98.11 99.06 97.76 98.87
X-GAN 99.41 99.71 98.66 99.33 99.21 99.60 99.42 99.71 99.19 99.59
X-GAN (MAX) 99.89 99.94 99.66 99.83 99.73 99.86 99.80 99.89 99.57 99.78
Table 1: Comparison X-GAN With other models on GSS-RetVein, OCTA-500 and ROSE. ($\times$100%)
Data Collection and Quality Control: our partner medical institutions’s institutional review board (IRB) approved this study, which followed the principles of the Declaration of Helsinki. Since the study was retrospective, the IRB waived the requirement for informed consent from patients.
4.1.2 OCTA-500
OCTA-500 [30] focuses on retinal blood vessel research, providing 6mm×$\times$6mm (300 images, ID range 10001 to 10300) and 3mm×$\times$3mm (200 images, ID range 10301 to 10500) fields of view for wide-field vessel analysis and high-resolution microvascular studies. It includes detailed annotations for arteries, veins, large vessels, capillary networks, and 2D/3D Foveal Avascular Zone, facilitating analysis of retinal vessels structures and pathological changes.
4.1.3 ROSE
ROSE [40] is an open-source collection designed for retinal blood vessel segmentation using OCTA images. It consists of two subsets: ROSE-1 includes 117 OCTA images from 39 subjects, covering a 3mm×$\times$3mm foveal-centered area with a resolution of 304$\times$304 pixels, providing both centerline-level and pixel-level vessel annotations. ROSE-2 contains 112 OCTA images from 112 eyes, focusing on the superficial vascular complex within a 3mm×$\times$3mm area, resized to 840$\times$840 pixels, with centerline-level annotations.
4.2 Implement
4.2.1 Hyperparameter
For X-GAN, its generator adopts a ResNet 9-block architecture, while the discriminator utilizes PatchGAN (70$\times$70). The optimizer is Adam with a learning rate of $2\times 10^{-4}$ and momentum parameters $\beta_{1}=0.5$, $\beta_{2}=0.999$. X-GAN is trained for 50 epochs, with a linear learning rate decay to 0 after 25 epochs, with evaluation performed using 10-fold cross-validation. Data augmentation includes: random rotations ($k\times 90^{\circ}\pm 10^{\circ}$), flipping and contrast adjustments.
The hardware specifications for training and testing include 2 Tesla V100 GPUs (2 $\times$ 32GB), 64GB of RAM, 8 CPU cores per node, and a total of 6 nodes.
4.2.2 Evaluation Metric
We use Intersection over Union (IoU), also known as the Jaccard Index, to measure similarity between the predicted segmentation and the ground truth: $\text{IoU}=|P\cap G|/|P\cup G|$. This metric ranges from 0 to 1, with 1 indicating perfect segmentation. Additionally, we use the Dice Coefficient, another metric for segmentation accuracy that calculates overlap between the prediction and ground truth: $\text{Dice}=2\times|P\cap G|/(|P|+|G|)$. In both formulas, $P$ represents predicted vessel pixels, $G$ the ground truth vessel pixels, $|P\cap G|$ the count of correctly predicted vessel pixels, and $|P\cup G|$ the combined set of ground truth and predicted vessel pixels. All in all, IoU is used to evaluate segmentation accuracy of the main vessels, while the Dice Coefficient assesses accuracy at finer vessel ends relative to the main vessels. For vascular structural enhancement, we use SSIM (Structural Similarity Index) and MSE (Mean Squared Error). SSIM (range: -1 to 1) assesses similarity based on luminance, contrast, and structure, while MSE quantifies pixel-wise differences, with lower values indicating higher similarity. While SSIM aligns with human perception, PSNR and MSE focus on absolute pixel differences.
(a) Comparison of models, trained on GSS-RetVein and OCTA and tested on ROSE.
(b) Comparison of models, trained on ROSE and tested on GSS-RetVein and OCTA.
(c) Comparison of models, trained on GSS-RetVein or OCTA and tested on each other.
Figure 6: Conducting mixed testing on different models using mixed datasets. ($\times$100%)
5 Experimental Result
Model Dataset
GSS-RetVein OCTA-500 ROSE
6mm 3mm 6mm 3mm 3mm
SSIM MSE SSIM MSE SSIM MSE SSIM MSE SSIM MSE
CycleGAN [66] 93.75 2.36 94.32 2.39 93.56 1.62 94.58 4.46 91.01 3.83
CycleGAN (ours) 95.21 1.88 95.89 1.85 94.82 1.73 96.04 3.10 95.32 2.16
Table 2: Comparison vessel structure refinement images with real OCTA images. ($\times$100%)
5.1 Comparative Experiment
As presented in Tab. 1, experimental findings demonstrate that while models such as MedSAM, U-Net++, and YOLOv11-x exhibit competitive performance, X-GAN consistently outperforms all competing models across varying dataset resolutions, achieving the highest scores and excelling in retinal vessel segmentation, where both of the evaluation matrixs are even close to 100%, highlighting its superior segmentation accuracy and underscoring its robustness in fine-grained vascular structures segmentation.
As shown in Fig. 6(a), given the comparable sizes of GSS-RetVein and OCTA-500 (retinal vessel subset) and the smaller size of ROSE, we trained models on the first two and tested on the latter. X-GAN consistently achieved the highest IoU and Dice scores on ROSE, outperforming S2VNet, CauSSL, and MedSAM. To further assess generalization, we trained models on ROSE and tested them on GSS-RetVein and OCTA-500 (Fig. 6(b)), where X-GAN again achieved the highest scores, followed by MedSAM and CauSSL. Finally, Fig. 6(c) evaluates models across different scanning resolutions, with X-GAN consistently leading in segmentation accuracy. While MedSAM and UniverSeg performed well, U-Net and its variants showed strong but relatively lower performance. These results highlight GSS-RetVein’s high quality and X-GAN’s superior segmentation capability.
Note: the highest indicator is shown in bold, while the second highest is marked with an underline. We do not consider X-GAN (MAX) in all Tabs. 1, 6 and 3.
5.2 Ablation Experiment
For baseline GAN model, GAN [13] and CycleGAN [66] are selected. For baseline segmentor, U-Net [49], U-Net++ [64], Attention U-Net [44] and MedSAM [39] are selected. For X-GAN (CycleGAN (ours) + DFS), the dirrerent values of $R_{min}$ are tested to evaluate the model segmentation accuracy. As shown in Tab. 3, since main blood vessels make up a small proportion, the value calculated is also relatively low. When the minimum radius ratio is found, like X-GAN (MAx), aligning with biostatistical properties [22, 12, 5, 43, 7]—segmentation accuracy peaks, reinforcing the model’s strong clinical interpretability.
Model Dataset
GSS-RetVein
6mm 3mm
IoU Dice IoU Dice
GAN [13] + U-Net [49] 92.89 96.31 92.44 96.07
GAN [13] + U-Net++ [64] 93.62 96.70 92.87 96.30
GAN [13] + Attention U-Net [44] 93.56 96.67 93.12 96.44
GAN [13] + MedSAM [39] 93.73 96.76 92.76 96.24
GAN [13] + DFS ($R_{min}$=0.2) 95.81 97.86 95.24 97.56
CycleGAN [66] + U-Net [49] 93.44 96.61 93.42 96.60
CycleGAN [66] + U-Net++ [64] 93.72 96.76 93.62 96.70
CycleGAN [66] + Attention U-Net [44] 93.79 96.80 93.45 96.61
CycleGAN [66] + MedSAM [39] 94.02 96.92 94.37 97.10
CycleGAN [66] + DFS ($R_{min}$=0.2) 96.23 98.08 96.42 98.18
CycleGAN (ours) + U-Net [49] 94.56 97.20 94.52 97.18
CycleGAN (ours) + U-Net++ [64] 94.98 97.43 94.36 97.10
CycleGAN (ours) + Attention U-Net [44] 95.06 97.47 95.12 97.50
CycleGAN (ours) + MedSAM [39] 95.33 97.61 95.21 97.55
X-GAN 99.41 99.71 98.66 99.33
X-GAN (MAX) 99.89 99.94 99.66 99.83
Table 3: Ablation experiment of X-GAN. ($\times$100%)
As shown in Fig. 7(a), the optimal choice of $R_{min}$ (0.2) enables X-GAN to significantly outperform other models in segmentation, which is so high that it is nearly perfect, approaching 100%. When $R_{min}=1$, X-GAN segments only the largest vessels, resulting in an extremely low segmentation index as only the starting points meet the threshold. When $R_{min}$ is 0, it is equivalent to no filtering, directly the original image.
(a) $R_{min}$
(b) The Number of Nodes $N$
Figure 7: Ablation experiments of key parameters. ($\times$100%)
Also, as shown in Eq. 1 and Fig. 7(b), a greater number of generated nodes $N$ leads to smoother vessel edges, a more complete vascular structure, and improved segmentation accuracy. Given that the R values of X-GAN and X-GAN (MAX) are relatively low, their practical significance indicates minimal differences in the segmented vessels, as the number of nodes remains similar. Consequently, their curves largely overlap, with X-GAN (MAX) exhibiting a slight increase.
5.3 Evaluation of Vessel Structure Refinement
The comparison between CycleGAN [66] and CycleGAN (ours revised version) shows a high structural similarity, with SSIM values ranging from 91.01% to 96.04%. CycleGAN (ours) consistently outperforms CycleGAN, achieving higher SSIM and lower MSE across all datasets. The small SSIM difference (less than 4.31%) indicates that both models generate highly similar images, but CycleGAN (ours) preserves image structure better and reduces pixel-level errors, making it the superior model.
Human Evaluation: the medical team verified the corrected images. Luckily, since this study only aimed to segment only the main blood vessels, the small size of the capillaries caused errors in radius generation, which affected the accuracy of the indicators.
Indicator SLSMs DFS (ours)
GPU $\checkmark$ $\times$
Annotation $\checkmark$ $\times$
Training $\checkmark$ $\times$
Explainability $\times$ $\checkmark$
2D and 3D $\times$ $\checkmark$
Generalization $\times$ -
Table 4: Comparison of performance indicators between DFS (our) and SLSMs. ($\times$100%) (- indicates strong performance for a specific class.)
(a) Segmentation of Main Retinal Vessels (2D)
(b) Original Image (2D)
(c) Original Image (3D)
(d) Segmentation (2D)
(e) Segmentation (3D)
Figure 8: Segmentation results of X-GAN on GSS-RetVein (2D and 3D).
5.4 Efficiency Evaluation
All in all, as shown in Tab. 4, compared to SLSMs, our segmentor offers these advantages: faster inference (DFS does not need GPU and training, but SLSMs need GPU for training and testing), eliminates the need for large-scale annotations, and ensures full preservation of primary vessel integrity. It features an adjustable capillary filtering parameter $R_{min}$, maintains strong generalization across all vessel OCTA datasets without retraining, and is directly applicable to 3D extensions without additional design.
What is more, in more detail, the advantages of our approach are as follows in Tab. 5, our segmentor has demonstrated unsurpassed superiority.
Model Details
Inference Time Parameters FLOPs GPU Utilization
U-Net [49] 21ms 31M 98 low
U-Net++ [64] 62ms 76M 304 high
Attention U-Net [44] 83ms 62M 215 slightly high
MedSAM [39] 425ms 109M 1250 extremely high
DFS (ours) 0.018ms 0M 0.00028 -
DFS (ours, GPU based) $\leq$ 0.01ms 0M 0.00028 extremely low
Table 5: Comparison of performance details between DFS (our) and SLSMs.
5.5 Visualization
As shown in Fig. 8, our 2D and 3D segmentation results exhibit smooth, well-defined vessel curves, confirming our prior analysis. In Fig. 8(a), the segmented vessels are exceptionally smooth, free of pixel artifacts or overflow. DFS effectively filters out vessels below the threshold, eliminating noise and misclassified capillaries. Extending this to 3D, we achieve precise main vessel extraction, with Fig. 8(e) preserving vessel integrity and offering a clearer structure than the original Fig. 8(c).
5.6 Analysis
Optimal selection of $R_{min}$ values allows X-GAN to achieve exceptionally high segmentation accuracy, outperforming the second-best model by nearly 3 percentage points, nearly 100% (Tabs. 1, 6 and 3). Unlike many SLSMs that have been magically tuned to have very poor interpretability, $R_{min}$ is based on biostatistics of glacucoma [22, 12, 5, 43, 7]. We only knew the biological information of the main blood vessel radius and the form of its data expression, and just adjusted X-GAN to achieve amazing segmentation results. Studies [34, 56] have also shown the similar idea: with a deep understanding of data characteristics, the right training strategies, and optimized parameters, even CNNs can outperform transformer-based models. There are many examples: Faster R-CNN [48] and Mask R-CNN [16], YOLOv3 [47], YOLOv4 [4] and YOLOv5 [17], DeepSeek [14, 33] and ChatGPT [1].
To the best of our knowledge, X-GAN is the best model currently, tailored to segment OCTA retinal main vessel.
6 Conclusion
In this paper, we propose X-GAN, an unsupervised model for ultra-high-precision segmentation of OCTA retinal main vessels, integrating biostatistical vessel radius properties with a GAN-enhanced DFS algorithm to achieve near-perfect accuracy without labeled data, training, or high-performance GPUs. We also introduce GSS-RetVein, a high-resolution dataset featuring both 2D and 3D retinal vascular structures, offering sharper vessel boundaries and detailed capillary networks compared to existing datasets, establishing it as a superior benchmark for vessel segmentation. Extensive experiments demonstrate that X-GAN consistently surpasses SOTA models, achieving nearly 100% segmentation accuracy, confirming its robustness and generalizability. Our DFS-based segmentation effectively isolates main vessels while filtering out capillary noise, ensuring strong biological interpretability and clinical relevance.
Acknowledgements
This research is sponsored by NASA 80NSSC22K0144, UTSW GMO No. 241213, and NIH 1R01AG083179-01.