### v4

# GlaGAN: A Generative Unsupervised Model for High-Precision Segmentation of Retinal Main Vessels toward Early Detection of Glaucoma

Abstract
Structural changes in the main retinal blood vessels are critical biomarkers for glaucoma onset and progression. Identifying these vessels is essential for vascular modeling yet highly challenging. This paper introduces GlaGAN, an unsupervised generative AI model for segmenting main blood vessels in Optical Coherence Tomography Angiography (OCTA) images. The process begins with the Space Colonization Algorithm (SCA) to rapidly generate vessel skeletons, including radius estimations. By synergistically integrating generative adversarial networks (GANs) with biostatistical modeling of vessel radii, GlaGAN efficiently reconstructs 2D and 3D representations, achieving nearly 100% segmentation accuracy without requiring labeled data or high-performance computing resources. To address data scarcity, we also present GSS-RetVein, a high-definition mixed 2D/3D glaucoma retinal dataset featuring clear capillary structures. Designed for robustness testing, GSS-RetVein incorporates controlled noise while maintaining sharp capillary boundaries in 2D and enhancing 3D vascular reconstruction for blood flow prediction and glaucoma progression simulations. Experimental results demonstrate GSS-RetVein outperforms existing datasets in evaluating main vessel segmentation. Code and dataset are available: https://github.com/VikiXie/SatMar8.
{IEEEkeywords}
Glaucoma, Generative Adversarial Network, Space Colonization Algorithm, Medical Imaging, Main Blood Vessel Extraction, Image Segmentation
1 Introduction
[…]
Glaucoma is a leading cause of irreversible blindness, often progressing silently until significant vision loss occurs [40, 37, 38, 42, 29]. Recent research suggests that choroidal microvasculature dropout may serve as a potential biomarker for disease progression [59, 60]. While Optical Coherence Tomography Angiography (OCTA) allows detailed visualization of these vascular changes, accurately quantifying choroidal vessel density remains a major challenge [1, 17].
Existing medical imaging analytics in glaucoma research predominantly depend on computer vision techniques [2, 46, 45, 44, 17, 37, 38, 36, 39, 41, 40, 47]. In recent years, supervised learning segmentation models (SLSMs) [10, 47, 12, 13, 45, 46, 44, 9, 8, 2, 3, 14, 7, 15] have become particularly prominent in this field. However, SLSMs may not be well-suited for detecting choroidal vessels in OCTA images. They rely heavily on labeled data, but manually annotating choroidal blood vessels in OCTA images is extremely challenging with current devices [1, 4]. OCTA images offer high-resolution, depth-resolved visualization of retinal and choroidal microvasculature without the need of contrast dye, making them more effective than traditional fundus imaging for detecting vascular abnormalities [1, 4]. However, segmenting choroidal vessels is difficult due to irregular patterns, intersecting pathways, dense capillary networks [43, 42], and complex capillary structures. Moreover, the presence of major blood vessels, which must be excluded, further complicates the labeling process [42, 29].
Glaucoma researchers have revealed that retinal blood vessels follow biostatistical relationships, meaning that vessel radius variations along branching structures [32, 17]. raising the question of whether segmentation can bypass pixel-based mapping and leverage vessel radius as a key marker. To address this, we propose GlaGAN, an unsupervised model that integrates GANs [34] with retinal vessel biostatistics for precise main vessel segmentation from OCTA images, eliminating the need for labeled data or extensive training. To tackle data scarcity, we introduce GSS-RetVein, a benchmark dataset with clear capillary distribution in 2D and 3D, enhancing segmentation, 3D reconstruction, and blood flow prediction.
In summary, the contributions of this paper are three-fold:
  • We introduce GlaGAN, a generative AI-powered unsupervised segmentation model tailored for extracting main blood vessels from OCTA images.
  • We propose the GSS-RetVein, a unique and pioneering dataset focused exclusively on retinal vascularity. GSS-RetVein consists of 2D and 3D imaging data, and also preserves intricate capillary details without compromising resolution, ensuring sharp blood vessel boundaries and minimal pixel-induced blurring.
  • Experimental results of GlaGAN on GSS-RetVein, OCTA-500 [1] and ROSE [4] for retinal main vessel segmentation have demonstrate that GlaGAN outperformed state-of-the-art (SOTA) cureent models, achieving nearly 100% segmentation accuracy. Also, experimental results of different models on GSS-RetVein have shown the superiority of GSS-RetVein.
2 Related Work
2.1 Datasets for Glaucoma
Many datasets of glaucoma focus on computer vision tasks [37, 38, 36, 39, 41, 40, 44, 4, 1, 54, 55, 53]. For inner retinal blood vessels, datasets like LAG [44], OCTA-500 [1] and ROSE [4], provide various-sizeimages (e.g., 3mm, 6mm). However, these images represent only a small portion of the total dataset. Some datasets [54, 55, 53, 37, 38, 36, 39, 41, 40, 44] offer fundus images, and the capillaries are nearly indistinguishable to the naked eye. There is a shortage of high-quality, intuitive data, such as detailed images of inner 3mm×$\times$3mm retinal vessels, and existing datasets often lack sufficient clarity. Consequently, few studies focus on structural changes in major retinal blood vessels in glaucoma.
Figure 1: The Architecture of GlaGAN
2.2 Computer Vision for Glaucoma
Even when data is accessible, challenges in labeling and the subjective variability among doctors may introduce biases or hinder optimal model fitting [37, 38, 36, 39, 41, 40, 47]. Generative AI can compensate for data deficiencies and labeling issues, though it often requires human verification [51] and generated data also face challenges in labeling. However, GAN [34, 35, 19, 25, 28, 26] is an unsupervised model which can solve both issues: data scarity and labeling. GAN-based segmentation methods are generally designed in conjunction with a dedicated segmentation network. For the segmentor part $S$, many pre-trained models like CNNs [46, 45, 44], U-Net [10, 7, 47, 12], ViT [23], Mask R-CNN [9] and MedSAM [8] can be used for final segmentation. Most recent studies from CVPR [18, 19, 21, 22], IEEE Transactions on Medical Imaging [17, 26, 27, 28] and other top conferences or journals [20, 24, 25] also follow this way.
2.3 Motivation
For GAN+SLSM [17, 26, 21, 27, 28, 19], or SLSMs [10, 47, 12, 13, 45, 46, 44, 9, 8, 2, 3, 14, 7, 15], pixel-level representations can be subjective, as different doctors may annotate the details of the same object differently [40, 1, 2]. In OCTA imaging, retinal blood vessels exhibit a uniform color change and have a shape closely resembling a path. They can be represented using coordinates along with road width. Space Colonization Algorithm (SCA) [50] just meets this requirement. Rather than generating pixel-wise vessel segmentations, we formulates the vascular network as a structured graph representation, ensuring topological consistency and robust connectivity. The graph consists of the following components, shown in Eq. 1:
$V=\left\{(x_{i},y_{i},z_{i}|r_{i})\right\}_{i=1}^{N}$ (1)
where ($x_{i}$, $y_{i}$, $z_{i}$) are the spatial coordinates of the vessel centerline points, $r_{i}$ represents the local vessel radius and $N$ means the number of nodes. This structured representation is initially generated using the SCA [50, 57, 58], which models vessel growth based on attraction points and bifurcation principles. However, the initial vessel maps exhibit domain discrepancies in contrast and noise characteristics compared to real OCTA images. To mitigate this gap, we employ a tuned GAN based on CycleGAN [35] as a structural refinement module, aligning vessel representations with real OCTA distributions while preserving vascular topology.
We introduce the Depth-First Search (DFS, a nineteenth century algorithm)-based segmentation approach [52], applying DFS directly to structured (coordinates, radius) data before image rendering. Unlike GAN+SLSM [17, 26, 21, 27, 28, 19], which segments generated images mixed with real data, our method intercepts the GAN generation process, ensuring segmentation aligns with biostatistical vessel radii. The DFS threshold, $R_{min}$, is defined as the minimum radius ratio of main vessels, following established biostatistical principles [29, 30, 31, 33, 32]. Also, we introduce GSS-RetVein, a dataset capturing rich details of both main vessels and capillaries. The dense capillaries of our dataset introduce noise, effectively testing model robustness for SLSMs. This detail aids vascular reconstruction, blood flow prediction, and glaucoma progression simulation by enhancing edge and endpoint learning [1, 2, 3]. Inspired by Harvard glaucoma dataset series [37, 38, 36, 39, 41, 40], we integrate fairness measures to minimize bias and ensure equitable model performance.
3 Methodology
The complete framework of GlaGAN is illustrated in Fig. 1 where red indicates that GlaGAN does not utilize these things in Segmentor $S$, whereas GAN+SLSMs do. and it consists of two parts: Vessel Structure Refinement Module Module and Segmentor.
3.1 Vessel Structure Refinement Module
Given a vessel structure $X_{s}$ via SCA, our generator $G$ learns to transform it into a realistic OCTA-like representation $X_{r}^{\prime}$, expressed as $G(X_{s})$, while $X_{r}^{\prime}$ undergoes style adaptation to align with the contrast and noise characteristics of real OCTA images $X_{r}$. To enhance structural consistency, we modify the original CycleGAN architecture by removing the inverse generator and incorporating a segmentation consistency loss via the segmentor $S$. This modification prevents CycleGAN from altering the vessel topology, ensuring that vessel segmentation remains biologically meaningful, as illustrated in Eq. 2:
$L_{\text{seg}}=\mathbb{E}_{X_{r}}\left[\|S(X_{r}^{\prime})-S(X_{r})\|_{1}\right]$ (2)
where $S(X)$ extracts vessel masks from OCTA images, enforcing structural consistency during transformation. This constraint ensures that CycleGAN adapts contrast while strictly preserving vessel topology. Consequently, our total loss function is formulated as follows:
$L_{\text{GAN}}$ $=\mathbb{E}_{X_{r}}[D(X_{r})]-\mathbb{E}_{X_{s}}[D(G(X_{s}))]$ (3)
$L_{\text{GP}}$ $=\lambda\mathbb{E}_{\hat{X}}\left[\left(\|\nabla_{\hat{X}}D(\hat{X})\|_{2}-1\right)^{2}\right]$ (4)
$L_{\text{total}}$ $=L_{\text{GAN}}+\lambda_{\text{GP}}L_{\text{GP}}+\lambda_{\text{seg}}L_{\text{seg}}$ (5)
where Eq. 3 is the Wasserstein Adversarial Loss, Eq. 4 is the Gradient Penalty for Stability ($\hat{X}$ is the reference vessel representation) and Eq. 5 is the Structural Consistency via Segmentation Loss. Upon training convergence, the generator produces a vessel structure that not only exhibits realistic contrast properties but also preserves anatomical coherence, ensuring fidelity to real OCTA characteristics.
3.2 Segmentor
Rather than utilizing pixel-wise segmentation maps, we directly extract primary vessels from the $(x_{i}^{G},y_{i}^{G},z_{i}^{G}|r_{i}^{G})$ representation using DFS based graph traversal, which effectively isolates large vessels while filtering out capillaries.
First, we construct a graph $G=(V,E)$, where nodes $V$ represent vessel centerline points and edges $E$ connect adjacent vessel points based on local vessel connectivity. A vessel segment $e_{ij}$ between nodes $v_{i}$ and $v_{j}$ (the direction is from $i$ to $j$) is assigned a weight, as shown in Eq. 6:
$w(e_{ij})=r_{i}$ (6)
where larger vessels are prioritized in the traversal process. To extract primary vessels, we apply radius thresholding and DFS traversal. For radius filtering, we define a minimum main vessel radius $R_{min}$ and retain only Eq. 7:
$V_{\text{main}}=\left\{(x_{i},y_{i},z_{i}|r_{i})\mid r_{i}\geq(R_{\min}*r_{max})\right\}$ (7)
For DFS traversal, it consists of three steps:
  • select the optic disc region as the root node
  • recursively traverse the largest connected component using DFS
  • stop when no further vessel segments satisfy the radius constraint
when all three steps finished, construct final extracted vessel structure $G_{main}$. This approach effectively retains only the primary vascular structure, removing small capillaries and noise. This method can be summarized as the following Algorithm 1.
Algorithm 1 Primary Vessel Extraction via DFS
0:  Vessel Structure $V=\{(x_{i},y_{i},z_{i}\mid r_{i})\}_{i=1}^{N}$
0:  Extracted Primary Vessel Graph $G_{\text{main}}$
1:  Step 1: Vessel Graph Construction
2:  for all vessel point $(x_{i},y_{i},z_{i}\mid r_{i})\in V$ do
3:     Identify Neighboring Vessel Points to Form Edge Set $E$
4:     Assign Edge Weight $w(e_{ij})=r_{i}$
5:  end for
6:  Step 2: Vessel Filtering
7:  Apply Radius Threshold $R_{\text{min}}$
8:  $V_{\text{main}}=\{(x_{i},y_{i},z_{i}\mid r_{i})\mid r_{i}\geq R_{\text{min}}\cdot r_{\max}\}$
9:  Remove Weakly Connected Components
10:  Step 3: DFS-Based Traversal
11:  Initialize DFS from Optic Disc Region
12:  for all point $(x_{i},y_{i},z_{i}\mid r_{i})\in V_{\text{main}}$ do
13:     Recursively Explore Vessel Network via DFS
14:     Preferentially Expand along Thicker Vessels
15:  end for
16:  Step 4: Output
17:  return  $G_{\text{main}}$
The key aspect of Algorithm 1 is the threshold $R_{min}$. Instead of the final image, we analyze intermediate outputs $S(X_{r})$, CSV data (Fig. 2(e)), and GAN-derived radius statistics (Fig. 2(d)). For a sample image (Fig. 2(a)), we computed a Mean (85.31) and Standard Deviation (88.67), determining $R_{min}$ as 0.1968 $r_{max}$. Across datasets, values were 0.2011 (GSS-RetVein), 0.2102 (OCTA-500), and 0.1982 (ROSE), standardized to 0.2 for consistency. Ophthalmology studies [29, 30, 31, 33, 32] confirm 18-26% capillary coverage in OCTA images, validating this threshold.
(a) GSS-RetVein
(b) 3D Visualization
(c) Statistical Distribution
(d) Radius Statistics
(e) Vascular Data
Figure 2: 3D Visualization of the Pixel Distribution of One Sample, and its Gaussian Statistical Distribution
4 Dataset and Implementation
4.1 Dataset
4.1.1 GSS-RetVein
GSS-RetVein is a unique and pioneering dataset focused exclusively on retinal vascularity. The dataset features an exceptionally clear distribution of capillaries. It comprises 550 images, both 2D and 3D, from 250 subjects, ensuring fair representation across gender and race. We collected retinal vascular data for both eyes (OS and OD) of each subject, with a gender split of 48.8% male and 51.2% female. Racial distribution reflects local glaucoma clinic demographics: 80.4% white, 14.6% Black, with Asians and other minority groups comprising the remainder. Additional data includes age (47.1 $\pm$ 24.8 years) and examination dates from 2023 to 2024.
GSS-RetVein’s original images come from the latest Intalight device [^1] https://intalight.com/, offering superior detail compared to OPTOVUE [^2] https://www.visionix.com/ (Fig. 3). As shown in Fig. 3(a) and Fig. 3(c), Intalight images feature clearer vessel boundaries and capillary distributions, enhancing segmentation accuracy. In supervised learning, their minimal pixel variation introduces valuable noise for testing model robustness [3]. For 3D reconstruction and evolution prediction, these high-resolution images aid in vascular modeling and blood flow analysis [2], making Intalight data ideal for both tasks.
(a) Intalight
(b) OPTOVUE
(c) Intalight
(d) OPTOVUE
Figure 3: Comparison of Data Collected by Different Devices for the Same Patient
For GSS-RetVein, we eliminate red and green baselines from Intalight device images through a multi-step process. First, we apply a dark channel algorithm [5] to remove light artifacts, followed by unsharp masking [6] to enhance high-frequency details. The final images (Fig. 4(a)) undergo manual verification for quality assurance. As shown in Fig. 4, while main vessels remain visible across datasets, magnification reveals that OCTA-500 [1] and ROSE [4] blur capillaries and vessel ends. Even the unprocessed baseline image (Fig. LABEL:cid) offers superior vessel clarity compared to other datasets (Fig. 4(c), Fig. 4(b)). Among them, Fig. 3(a) and Fig. 3(b) are images of deep angio, and Fig. 3(c) and Fig. 3(d) are images of superficial angio. They all are images of inner retina vessel.
(a) GSS-RetVein
(b) OCTA-500
(c) ROSE
Figure 4: Comparison of GSS-RetVein with Other Datasets
For the representation of 3D imaging of GSS-RetVein, as shown in Fig. 5, capillaries stand out from the thicker main blood vessels by attaching to them and narrowing at their ends (Fig. 5(c)). The elaborate architecture of blood vessels and the intertwining capillaries contribute to an extremely intricate vascular layout. 3D imaging part of GSS-RetVein is formatted as a JSON file, which includes 3D X, Y, and Z coordinates and the corresponding radius pixel distribution values. For annotation (2D, image format: PNG), our team includes two glaucoma specialists who handle labeling, verification, and manual evaluation, utilizing the Labelme [^3] https://github.com/wkentaro/labelme.
(a) Top View
(b) Front View
(c) Detail
Figure 5: 3D Imaging of Retinal Vessel from Various Angles
In general, GSS-RetVein is a dataset that combines 2D and 3D images with extremely high image resolution, and is particularly dedicated to retinal vascular segmentation.
Table 1: Comparison GlaGAN With Other Models on GSS-RetVein, OCTA-500 and ROSE ($\times$100%)
Model Dataset
GSS-RetVein OCTA-500 ROSE
6mm 3mm 6mm 3mm 3mm
IoU Dice IoU Dice IoU Dice IoU Dice IoU Dice
U-Net 95.14 97.51 95.23 97.56 96.93 98.44 98.75 99.37 97.94 98.96
U-Net++ 95.68 97.81 96.64 98.29 97.65 98.81 97.96 98.97 97.52 98.75
Attention U-Net 95.78 97.84 96.04 97.98 97.05 98.50 98.61 99.30 97.92 98.95
Mask R-CNN 92.35 96.02 91.35 95.48 92.06 95.87 93.50 96.64 91.23 95.41
YOLOv11-x 95.81 97.86 96.55 98.24 95.93 97.92 97.66 98.82 97.83 98.89
MedSAM 96.38 98.16 96.47 98.20 96.01 97.95 98.94 99.47 97.89 98.93
CauSSL 95.23 97.56 95.47 97.68 96.32 98.04 98.15 99.08 97.80 98.89
UniverSeg 95.30 97.58 95.52 97.70 96.58 98.17 98.42 99.22 97.07 98.01
S2VNet 95.68 97.81 95.36 97.63 96.45 98.11 98.27 99.15 97.92 98.95
Tyche 95.80 97.85 95.19 97.53 96.28 98.02 98.11 99.06 97.76 98.87
GlaGAN 99.41 99.71 98.66 99.33 99.21 99.60 99.42 99.71 99.19 99.59
Table 2: Conducting Mixed Validation on Different Models using Mixed Datasets ($\times$100%)
Model Dataset
ROSE GSS-RetVein OCTA-500 GSS-RetVein (Test) OCTA-500 (Test)
GSS-RetVein OCTA-500 ROSE OCTA-500 (Train) GSS-RetVein (Train)
3mm 6mm 3mm 6mm 3mm
IoU Dice IoU Dice IoU Dice IoU Dice IoU Dice IoU Dice IoU Dice IoU Dice
U-Net 98.34 99.16 98.32 99.15 93.89 96.85 94.23 97.03 96.21 98.07 95.23 97.56 96.98 98.47 96.81 98.38
U-Net++ 98.99 99.49 98.78 99.39 94.08 96.95 94.46 97.15 95.73 97.82 97.01 98.51 97.59 98.78 96.93 98.44
Attention U-Net 98.54 99.26 98.23 99.11 94.07 96.94 94.52 97.18 96.99 98.47 96.09 98.01 97.21 98.59 96.78 98.36
Mask R-CNN 96.37 98.15 96.31 98.12 90.73 95.14 91.89 95.77 94.23 97.03 94.11 96.97 96.68 98.31 95.96 97.94
YOLOv11-x 98.88 99.44 98.72 99.36 94.11 96.97 94.49 97.17 97.32 98.64 96.54 98.24 96.12 98.02 96.63 98.29
MedSAM 99.04 99.52 98.98 99.49 95.01 97.41 95.42 97.66 97.54 98.75 95.87 97.89 98.23 99.11 97.46 98.72
CauSSL 96.83 98.38 95.89 97.90 96.98 98.47 97.83 98.88 99.12 99.56 98.81 99.40 94.87 97.33 95.34 97.62
UniverSeg 98.96 99.48 98.56 99.28 94.23 97.03 94.59 97.24 96.81 98.38 96.32 98.16 97.91 98.96 98.12 99.05
S2VNet 99.23 99.61 98.79 99.39 93.89 96.85 95.22 97.55 96.95 98.44 95.83 97.87 97.87 98.95 97.87 98.95
Tyche 98.78 99.38 98.43 99.21 94.22 97.02 94.98 97.46 96.88 98.41 96.11 98.02 97.84 98.94 97.68 98.87
GlaGAN 99.55 99.78 99.21 99.60 96.98 98.46 97.69 98.83 98.87 99.43 98.79 99.39 99.58 99.79 99.41 99.70
Data Collection and Quality Control: our partner medical institutions’s institutional review board (IRB) approved this study, which followed the principles of the Declaration of Helsinki. Since the study was retrospective, the IRB waived the requirement for informed consent from patients.
4.1.2 OCTA-500
OCTA-500 [1] focuses on retinal blood vessel research, providing 6mm×$\times$6mm (300 images, ID range 10001 to 10300) and 3mm×$\times$3mm (200 images, ID range 10301 to 10500) fields of view for wide-field vessel analysis and high-resolution microvascular studies. It includes detailed annotations for arteries, veins, large vessels, capillary networks, and 2D/3D Foveal Avascular Zone, facilitating analysis of retinal vessels structures and pathological changes.
4.1.3 ROSE
ROSE [4] is an open-source collection designed for retinal blood vessel segmentation using OCTA images. It consists of two subsets: ROSE-1 includes 117 OCTA images from 39 subjects, covering a 3mm×$\times$3mm foveal-centered area with a resolution of 304$\times$304 pixels, providing both centerline-level and pixel-level vessel annotations. ROSE-2 contains 112 OCTA images from 112 eyes, focusing on the superficial vascular complex within a 3mm×$\times$3mm area, resized to 840$\times$840 pixels, with centerline-level annotations.
4.2 Implement
Table 3: Comparison Vessel Structure Refinement Images with Real OCTA Images ($\times$100%)
Model Dataset
GSS-RetVein OCTA-500 ROSE
6mm 3mm 6mm 3mm 3mm
SSIM MSE SSIM MSE SSIM MSE SSIM MSE SSIM MSE
GAN 91.76 5.69 90.39 4.43 89.69 9.85 91.06 7.36 87.84 8.73
CycleGAN 93.75 2.36 94.32 2.39 93.56 1.62 94.58 4.46 91.01 3.83
GAN 93.62 3.29 93.87 3.79 91.28 6.98 93.19 5.54 91.73 4.18
CycleGAN 95.21 1.88 95.89 1.85 94.82 1.73 96.04 3.10 95.32 2.16
4.2.1 Hyperparameter
for GlaGAN, its generator adopts a ResNet 9-block architecture, while the discriminator utilizes PatchGAN (70$\times$70). The optimizer is Adam with a learning rate of $2\times 10^{-4}$ and momentum parameters $\beta_{1}=0.5$, $\beta_{2}=0.999$. GlaGAN is trained for 50 epochs, with a linear learning rate decay to 0 after 25 epochs, with evaluation performed using 10-fold cross-validation. Data augmentation includes: random rotations ($k\times 90^{\circ}\pm 10^{\circ}$), flipping and contrast adjustments.
4.2.2 Hardware
the hardware specifications for training and testing include 2 Tesla V100 GPUs (2 $\times$ 32GB), 64GB of RAM, 8 CPU cores per node, and a total of 6 nodes.
4.2.3 Evaluation Metric
we use Intersection over Union (IoU), also known as the Jaccard Index, to measure similarity between the predicted segmentation and the ground truth. This metric ranges from 0 to 1, with 1 indicating perfect segmentation. Additionally, we use the Dice Coefficient, another metric for segmentation accuracy that calculates overlap between the prediction and ground truth. All in all, IoU is used to evaluate segmentation accuracy of the main vessels, while the Dice Coefficient assesses accuracy at finer vessel ends relative to the main vessels. For vascular structural enhancement, we use Structural Similarity Index (SSIM) and Mean Squared Error (MSE). SSIM (range: -1 to 1) assesses similarity based on luminance, contrast, and structure, while MSE quantifies pixel-wise differences, with lower values indicating higher similarity. While SSIM aligns with human perception, MSE focus on absolute pixel differences.
Table 4: Ablation Experiment of GlaGAN ($\times$100%)
Model Dataset
GSS-RetVein OCTA-500 ROSE
6mm 3mm 6mm 3mm 3mm
IoU Dice IoU Dice IoU Dice IoU Dice IoU Dice
GAN U-Net 92.89 96.31 92.44 96.07 93.31 96.56 92.84 96.31 93.37 96.59
GAN U-Net++ 93.62 96.70 92.87 96.30 93.72 96.73 93.03 96.36 92.84 96.31
GAN Attention U-Net 93.56 96.67 93.12 96.44 93.65 96.70 92.85 96.31 93.42 96.61
GAN MedSAM 93.73 96.76 92.76 96.24 92.83 96.30 92.47 96.10 93.76 96.75
GAN DFS 95.81 97.86 95.24 97.56 94.83 97.33 94.87 97.35 94.73 97.27
CycleGAN U-Net 93.44 96.61 93.42 96.60 94.73 97.26 93.97 96.89 93.76 96.75
CycleGAN U-Net++ 93.72 96.76 93.62 96.70 94.58 97.18 94.36 97.05 93.19 96.46
CycleGAN Attention U-Net 93.79 96.80 93.45 96.61 94.59 97.18 94.29 97.01 94.03 96.89
CycleGAN MedSAM 94.02 96.92 94.37 97.10 93.83 96.72 94.37 97.19 94.18 96.96
CycleGAN DFS 96.23 98.08 96.42 98.18 96.68 98.33 96.27 98.13 95.83 97.84
CycleGAN U-Net 94.56 97.20 94.52 97.18 94.86 97.31 94.79 97.27 94.19 96.94
CycleGAN U-Net++ 94.98 97.43 94.36 97.10 94.83 97.29 95.09 97.41 94.27 96.98
CycleGAN Attention U-Net 95.06 97.47 95.12 97.50 95.07 97.39 94.97 97.33 94.39 97.04
CycleGAN MedSAM 95.33 97.61 95.21 97.55 95.36 97.55 95.11 97.41 94.82 97.27
GlaGAN 99.41 99.71 98.66 99.33 99.21 99.60 99.42 99.71 99.19 99.59
5 Experimental Result
5.1 Experimental Setup
We selected many SOTA models for experimental verification, which are: U-Net [10], U-Net++ [7], Attention U-Net [12], Mask R-CNN [9], YOLOv11-x [11], MedSAM [8], CauSSL [13], UniverSeg [14], S2VNet [15], Tyche [16], GAN [34] and CycleGAN [35]. Among them, we used tool Optuna [^4] https://optuna.org/ to optimize their parameters. All final experimental results are the average of ten cross-validation times. * means models’ hyperparameters have been optimized by us.
5.2 Comparative Experiment
As shown in Table 1, experimental results indicate that while MedSAM, U-Net++, and YOLOv11-x perform competitively, GlaGAN consistently outperforms them across resolutions, achieving near-perfect evaluation metrics and excelling in fine-grained retinal vessel segmentation. As shown in Table 2, the second row is the test data set, and the third row is the training data set. Models trained on GSS-RetVein and OCTA-500 were tested on ROSE, where GlaGAN achieved the highest IoU and Dice scores, outperforming S2VNet, CauSSL, and MedSAM. For generalization, training on ROSE and testing on GSS-RetVein and OCTA-500 confirmed GlaGAN’s superiority, followed by MedSAM and CauSSL. Across different scanning resolutions, GlaGAN consistently led in segmentation accuracy, with MedSAM and UniverSeg performing well, while U-Net variants lagged. These results affirm GSS-RetVein’s quality and GlaGAN’s segmentation excellence.
5.3 Ablation Experiment
For baseline GAN model, GAN and CycleGAN are selected. For baseline segmentor, U-Net, U-Net++, Attention U-Net and MedSAM are selected. For GlaGAN (CycleGAN (ours) + DFS), the dirrerent values of $R_{min}$ are tested to evaluate the model segmentation accuracy. As shown in Table 4, since main blood vessels make up a small proportion, the value calculated is also relatively low. When the minimum radius ratio is found, like GlaGAN, aligning with biostatistical properties [29, 30, 31, 33, 32], segmentation accuracy peaks, reinforcing the model’s strong clinical interpretability.
Table 5: Ablation Experiment of Hyperparameters ($\times$100%)
Hyperparameter Dataset
GSS-RetVein OCTA-500 ROSE
6mm 3mm 6mm 3mm 3mm
IoU Dice IoU Dice IoU Dice IoU Dice IoU Dice
Rmin 0.00 train 1.03 2.04 1.00 1.98 1.31 2.59 1.22 2.41 1.13 2.23
test 0.75 1.49 0.87 1.73 0.91 1.80 0.77 1.52 0.93 1.84
0.10 train 41.32 58.47 41.98 59.12 43.09 60.23 45.57 62.59 44.29 61.38
test 40.29 57.42 42.75 59.89 42.31 59.45 45.06 62.13 43.97 61.08
0.20 train 98.71 99.35 98.07 99.03 98.86 99.43 98.14 99.06 98.05 99.02
test 99.41 99.71 98.66 99.33 99.21 99.60 99.42 99.71 99.19 99.59
0.30 train 87.15 93.13 86.29 92.64 86.04 92.49 85.74 92.30 84.37 91.52
test 84.17 91.40 83.96 91.27 84.09 91.35 84.02 91.31 83.98 91.28
0.40 train 57.92 73.34 56.44 72.15 55.86 71.65 55.45 71.28 57.09 72.69
test 58.73 73.98 57.85 73.29 57.29 72.85 56.93 72.53 57.24 72.80
0.50 train 55.72 71.56 56.98 72.60 54.79 70.78 56.93 72.53 57.04 72.65
test 54.17 70.25 53.83 69.99 52.62 68.97 53.71 69.89 53.83 69.99
0.60 train 40.26 57.40 41.27 58.42 40.39 57.53 41.03 58.17 39.27 56.39
test 35.62 52.51 38.92 56.03 37.46 54.50 37.08 54.10 35.74 52.63
0.70 train 35.36 52.25 34.72 51.55 33.91 50.64 37.06 54.08 35.94 52.85
test 32.24 48.76 31.83 48.27 32.64 49.21 33.39 50.06 34.07 50.83
0.80 train 17.83 30.26 19.25 32.29 18.83 31.69 17.79 30.21 16.68 28.59
test 26.37 41.73 25.73 40.93 25.49 40.61 25.13 40.15 24.80 39.73
0.90 train 9.85 17.94 10.32 18.71 9.89 17.99 11.27 20.26 9.91 18.03
test 12.24 21.82 13.06 23.10 12.94 22.92 11.98 21.39 10.57 19.12
1.00 train 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
test 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Node 0 train 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
test 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
5000 train 6.34 11.92 7.81 14.49 6.28 11.82 6.43 12.09 7.02 13.12
test 6.06 11.43 5.86 11.07 5.93 11.19 6.18 11.64 6.58 12.35
10000 train 17.82 30.24 18.92 31.81 17.96 30.44 19.06 32.02 17.48 29.75
test 20.59 34.15 21.76 35.73 22.84 37.19 21.96 36.00 20.98 34.69
15000 train 20.58 34.14 21.96 36.00 20.84 34.50 21.37 35.21 20.71 34.29
test 28.73 44.63 29.64 45.71 30.74 47.04 28.93 44.87 29.74 45.83
20000 train 41.47 58.61 42.34 59.49 41.65 58.80 42.07 59.18 44.03 61.14
test 42.06 59.17 41.11 58.26 41.93 59.04 40.62 57.77 41.98 59.12
25000 train 49.38 66.12 52.14 68.53 50.39 66.96 52.97 69.26 51.07 67.60
test 52.01 68.41 51.49 67.98 51.76 68.21 52.44 68.79 53.06 69.35
30000 train 59.23 74.39 59.17 74.34 58.48 73.78 57.99 73.40 58.49 73.79
test 61.25 75.97 60.84 75.64 59.73 74.78 61.79 76.37 60.83 75.63
35000 train 62.33 76.79 63.34 77.56 62.15 76.64 61.96 76.49 62.15 76.64
test 62.51 76.96 64.08 78.12 63.24 77.47 62.44 76.89 63.19 77.43
40000 train 62.74 77.13 64.27 78.25 63.74 77.88 63.96 78.06 63.49 77.64
test 62.59 77.03 64.13 78.16 63.96 78.06 61.84 76.39 63.97 78.07
45000 train 63.37 77.57 65.03 78.81 65.17 79.57 65.03 78.81 65.93 79.41
test 65.72 79.30 66.34 79.77 64.44 78.39 65.39 78.95 66.67 80.00
50000 train 66.53 79.89 66.02 79.53 65.52 79.14 65.74 79.31 66.03 79.54
test 69.34 81.89 68.18 81.06 69.01 81.65 66.95 80.19 68.39 81.24
50000 train 71.24 83.20 72.47 84.02 69.08 81.71 68.79 81.52 69.28 81.83
test 73.69 84.85 72.61 84.12 74.17 85.15 75.36 85.94 74.13 85.10
55000 train 75.48 85.99 76.24 86.54 76.14 86.47 75.83 86.25 74.19 85.17
test 76.02 86.38 75.47 85.98 74.45 85.34 73.99 85.05 74.73 85.57
60000 train 79.28 88.47 78.96 88.23 79.48 88.61 77.20 87.13 75.96 86.35
test 78.49 87.94 77.85 87.52 77.02 87.03 79.36 88.52 75.89 86.29
65000 train 85.27 92.05 86.09 92.52 85.78 92.29 85.96 92.42 86.06 92.50
test 84.18 91.40 84.07 91.33 83.33 90.89 82.39 90.37 83.09 90.76
70000 train 97.46 98.71 97.28 98.62 97.08 98.52 96.96 98.46 97.07 98.51
test 98.32 99.15 98.29 99.14 99.03 99.51 99.01 99.50 98.76 99.37
…… train 98.71 99.35 98.07 99.03 98.86 99.43 98.14 99.06 98.05 99.02
test 99.41 99.71 98.66 99.33 99.21 99.60 99.42 99.71 99.19 99.59
As shown in Table 5, the optimal choice of $R_{min}$ (0.2) enables GlaGAN to significantly outperform other models in segmentation, which is so high that it is nearly perfect, approaching 100%. When $R_{min}=1$, GlaGAN segments only the largest vessels, resulting in an extremely low segmentation index as only the starting points meet the threshold. When $R_{min}$ is 0, it is equivalent to no filtering, directly the original image.
Also, as shown in Fig. ref1 and Table 5, a greater number of generated nodes $N$ leads to smoother vessel edges, a more complete vascular structure, and improved segmentation accuracy. Given that the R values of GlaGAN are relatively low, their practical significance indicates minimal differences in the segmented vessels, as the number of nodes remains similar. Consequently, their curves largely overlap, with GlaGAN exhibiting a slight increase.
5.4 Evaluation of Vessel Structure Refinement
As shown in Table 3, comparison between CycleGAN [35] and CycleGAN (ours revised version) shows a high structural similarity, with SSIM values ranging from 91.01% to 96.04%. CycleGAN (ours) consistently outperforms CycleGAN, achieving higher SSIM and lower MSE across all datasets. The small SSIM difference (less than 4.31%) indicates that both models generate highly similar images, but CycleGAN (ours) preserves image structure better and reduces pixel-level errors, making it the superior model.
5.5 Efficiency Evaluation
As shown in Table 6, compared to SLSMs, our segmentor offers these advantages: faster inference (DFS does not need GPU and training, but SLSMs need GPU for training and testing), eliminates the need for large-scale annotations, and ensures full preservation of primary vessel integrity. It features an adjustable capillary filtering parameter $R_{min}$, maintains strong generalization across all vessel OCTA datasets without retraining, and is directly applicable to 3D extensions without additional design.
Table 6: Comparison of Performance Indicators between DFS and SLSMs ($\times$100%)
Indicator SLSMs DFS (ours)
GPU $\checkmark$ $\times$
Annotation $\checkmark$ $\times$
Training $\checkmark$ $\times$
Explainability $\times$ $\checkmark$
2D and 3D $\times$ $\checkmark$
Generalization $\times$ -
Model Details
Inference Time Parameters FLOPs GPU Utilization
U-Net 21ms 31M 98 low
U-Net++ 62ms 76M 304 high
Attention U-Net 83ms 62M 215 slightly high
MedSAM 425ms 109M 1250 extremely high
DFS (CPU) 0.018ms 0M 0.00028 -
DFS (GPU) $\leq$ 0.01ms 0M 0.00028 extremely low
Table 7: Comparison of Performance Details between DFS and SLSMs
Compared to the segmentor, our adjusted DFS offers unmatched advantages, as shown in Table 7. It requires no separate training or labeling, as DFS is a parameter-efficient algorithm. By eliminating manual annotation, it removes subjective bias from doctors. Additionally, its radius-based segmentation avoids pixel mapping errors in SLSMs, such as interference at vessel edges, inaccuracies at junctions, and mask coverage errors. What is more, the key issue is that both SLSMs and GAN+SLSMs frameworks are susceptible to fitting problems in segmentation tasks. This arises from the training limitations of the SLSMs model and the nature of the data. In contrast, DFS, as an algorithm rather than a model, is unaffected by this issue.
5.6 Visualization
As shown in Fig. 6, 2D and 3D segmentation results exhibit smooth, well-defined vessel curves, confirming our prior analysis. In Fig. 6(a), the segmented vessels are exceptionally smooth, free of pixel artifacts or overflow. DFS effectively filters out vessels below the threshold, eliminating noise and misclassified capillaries. Extending this to 3D, we achieve precise main vessel extraction, with Fig. 6(e) preserving vessel integrity and offering a clearer structure than the original image (Fig. 6(c)).
(a) Segmentation of Main Retinal Vessels (2D)
(b) Original Image (2D)
(c) Original Image (3D)
(d) Segmentation (2D)
(e) Segmentation (3D)
Figure 6: Segmentation Results of GlaGAN on GSS-RetVein (2D and 3D)
5.7 Analysis
Optimal selection of $R_{min}$ values allows GlaGAN to achieve exceptionally high segmentation accuracy, outperforming the second-best model by nearly 3 percentage points, nearly 100% (Table 1, Table 2, Table 4). Unlike many SLSMs that have been magically tuned to have very poor interpretability, $R_{min}$ is based on biostatistics of glacucoma [29, 30, 31, 33, 32]. We only knew the biological information of the main blood vessel radius and the form of its data expression, and just adjusted GlaGAN to achieve amazing segmentation results. Unlike GAN+SLSMs or SLSMs, GlaGAN does not require additional data to avoid fitting issues. Its segmentor is efficient, computationally friendly, and free from pixel mapping errors associated with SLSMs segmentation.
To the best of our knowledge, GlaGAN is the best model currently, tailored to segment OCTA retinal main vessel.
6 Conclusion
In this paper, we propose GlaGAN, an unsupervised model for ultra-high-precision segmentation of OCTA retinal main vessels, integrating biostatistical vessel radius properties with a GAN-enhanced DFS algorithm to achieve near-perfect accuracy without labeled data, training, or high-performance GPUs. We also introduce GSS-RetVein, a high-resolution dataset featuring both 2D and 3D retinal vascular structures, offering sharper vessel boundaries and detailed capillary networks compared to existing datasets, establishing it as a superior benchmark for vessel segmentation. Extensive experiments demonstrate that GlaGAN consistently surpasses SOTA models, achieving nearly 100% segmentation accuracy, confirming its robustness and generalizability. Our DFS-based segmentation effectively isolates main vessels while filtering out capillary noise, ensuring strong biological interpretability and clinical relevance.
Acknowledgment
This research is partially supported by NASA under grant 80NSSC22K0144, UTSW GMO No. 241213, and NIH grant 1R01AG083179-01.